### v1

Abstract
Reservoir is an efficient network for time series processing. It is well known that network structure is one of the determinants of its performance. However, the topology structure of reservoirs, as well as their performance, is hard to analyzed, due to the lack of suitable mathematical tools. In this paper, we study the topology structure of reservoirs using persistent GLMY homology theory, and develop a method to improve its performance. Specifically, it is found that the reservoir performance is closely related to the one-dimensional GLMY homology groups. Then, we develop a reservoir structure optimization method by modifying the minimal representative cycles of one-dimensional GLMY homology groups. Finally, by experiments, it is validated that the performance of reservoirs is jointly influenced by the reservoir structure and the periodicity of the dataset.
keywords:
[…]
1 Introduction
Neural networks, inspired by the biological brain, have become a cornerstone in artificial intelligence and machine learning, enabling solutions to complex tasks in image recognition, natural language processing, and predictive analytics [1, 2]. Their structure—including layer count, neuron connectivity, and activation function type—directly determines functional capabilities, with early studies establishing foundational concepts like weight roles and nonlinearity from activation functions [3]. Moreover, a recent study has proposed the weight-agnostic neural networks [4], which can finish learning tasks without weight training but only with structure optimization. This shows that the structure of neural networks plays an essential role. However, as real-world problem complexity grows, a deeper understanding of task-optimized neural network structure design is needed.
The continuous improvement and optimization of neural network structures hold profound and far-reaching significance in advancing the practical application and theoretical development of artificial intelligence. From a performance perspective, a well-optimized structure can significantly enhance the network’s ability to extract features and learn patterns, enabling it to achieve higher accuracy and robustness in handling complex tasks. In terms of resource efficiency, optimizing the number of layers, neuron density, and connectivity patterns can effectively reduce the model’s parameter scale and computational complexity. Additionally, structural optimization plays a crucial role in enhancing the adaptability and interpretability of neural networks. By adjusting the structure to match the characteristics of specific tasks, the network’s generalization ability across different domains can be improved. Meanwhile, rational structural design can reduce the “black-box” nature of neural networks—for instance, through modular structural design, researchers can more clearly trace the flow of information and the contribution of each module to the final output, laying the foundation for interpreting the internal working mechanism of the network [5]. Moreover, recent studies indicate that the topology of a network’s structure significantly impacts its performance, suggesting that understanding and designing network architectures from a topological perspective is a crucial and effective approach [6].
Reservoir Computing (RC), as a distinctive paradigm within the machine learning landscape, has rapidly gained traction and widespread adoption owing to its inherent advantages that address key challenges in traditional neural network deployment. Specifically, RC exhibits a compact model architecture, which minimizes the demand for storage resources; enables rapid training processes by focusing optimization only on the readout layer rather than the entire network, significantly reducing computational overhead; delivers high prediction accuracy across various time-series and nonlinear tasks; and maintains excellent generalization performance when faced with unseen or noisy data—these attributes collectively make RC a preferred choice for applications ranging from real-time signal processing to complex system prediction [7, 8]. Nevertheless, a critical limitation of RC lies in its “black-box” nature, primarily stemming from the lack of clear interpretability regarding the internal dynamics of the reservoir. The reservoir, typically composed of a large number of randomly connected neurons, operates through complex nonlinear interactions, making it difficult to establish a direct link between its structural characteristics (e.g., connectivity density, neuron type, and topological arrangement) and its prediction performance. Consequently, investigating how the internal structure of the reservoir influences its prediction ability has become an increasingly attractive and essential research direction, as resolving this issue could not only unlock the potential to further optimize RC performance but also enhance its reliability and applicability in critical domains where interpretability is essential [9].
In practice, a reservoir is usually regarded as a digraph. In previous studies, the lack of mathematical tools to study the topology of digraphs made it difficult to topologically optimize the structure of the reservoir from an algebraic topological perspective. Now with the rapid development of GLMY homology theory, there are new effective tools for topology optimization of digraphs. The GLMY homology, also known as path homology and initially proposed by [10], represents a significant development in computational topology. As complex systems—ranging from social networks and neural connections to transportation and biochemical pathways—are increasingly modeled using directed graphs (digraphs), traditional homology theories, which are designed for undirected topological spaces, prove inadequate. They fail to capture crucial directional information, such as information flow, influence propagation, or causal relationships. GLMY homology overcomes this fundamental limitation by constructing a chain complex specifically based on directed paths within digraphs. Its distinct sensitivity to edge directions allows it to capture the flow and hierarchical organization of these paths, revealing hidden structural features and topological properties that are otherwise invisible to standard methods. In real-world digraphs, edge orientation often conveys critical information, such as influence direction in a social network or material flow in a manufacturing process. The directional awareness of GLMY homology provides a more accurate and nuanced representation of the digraph’s underlying topology. Persistent GLMY homology [11], a further development of GLMY homology, enhances its analytical capabilities by tracking how directed path structures evolve as parameters, such as edge weights, change. Through persistent GLMY homology, researchers can effectively capture structural features encoded in both edge directions and weights, making it a powerful tool for analyzing dynamic and weighted complex systems. This framework provides a new lens through which to analyze the structural properties of complex networks, particularly those where directionality is a defining feature.
Therefore, in this study, since a reservoir can be represented as a digraph, we present a novel method for reservoir optimization using GLMY homology. Firstly, we show that adding rings  [^1] While some literature uses the term “cycles”, we adopt “rings” to avoid conceptual conflicts with homology theory. [12, 13] in the reservoir digraph can improve the orthogonality of the reservoir adjacent matrix, hence can enhance the reservoir’s predictive capacity. This serves as the fundamental motivation for our approach. Our strategy focuses on strategically modifying the edge directions within the digraph to deliberately increase the number of rings, thereby optimizing the reservoir’s performance. Moreover, since these rings are a special type of representative cycle of one-dimensional GLMY homology, we can achieve our target by considering a subset of representative cycle of 1-d GLMY homology, i.e., the subset consists of all rings. We also validated by experiments that the prediction ability of a reservoir is related to both its structure and the topological features of the dataset.
In summary, the main contributions of our study are as follows:
  • Adding rings inside the reservoir digraph can improve the reservoir adjacent matrix’s orthogonality, which is equivalent to increase a certain type of representative cycle of GLMY homology.
  • We optimize reservoirs’ structure by expanding a certain subset of representative cycle (rings) of 1-d GLMY homology.
  • It is validated by experiments that the performance of reservoirs is related to both their structure and the characteristics of the dataset.
The remainder of this paper is organized as follows: We provide an overview of reservoir computing, reservoir optimization and GLMY homology in Section 2. Then, we present the key tools and the proposed method in Section 3 and 4. In Section 5, we show our experimental results and provide some discussions. Section 6 presents the conclusions of this study.
Figure 1: Illustration of the flowchart of our method and experimental procedures.
2 Related Work
In this section, we provide some related work about optimizing reservoir structure, and the GLMY homology and some of its applications.
2.1 Reservoir computing and reservoir optimization
Reservoir computing (RC) has emerged as a powerful and widely adopted machine learning paradigm over the past few decades, gaining considerable attention for its effectiveness in handling complex time-dependent data. Its application has extended across various domains, including time series prediction, nonlinear system identification [14], and speech recognition [15]. The core strength of RC lies in its architectural simplicity: a fixed, randomly generated recurrent neural network (the reservoir) projects the input into a high-dimensional feature space, while only a simple linear readout layer is trained. This separation of the nonlinear dynamics of the reservoir from the linear readout training process is a key reason for its exceptional computational efficiency and rapid training speed. This efficiency has led to remarkable results, such as those demonstrated by Jaeger et al. [16], who applied RC to chaotic time series prediction and achieved computational accuracy two orders of magnitude higher than traditional approaches. Beyond its core architecture, variations have further expanded its capabilities. For instance, Shougat et al. [17] employed Hopf oscillators as an intermediate reservoir, successfully demonstrating robustness to noise and preservation of predictive accuracy even with reduced training costs. In the financial sector, Wang et al. [18] utilized an RC network for forecasting stock market prices, where comparative experiments against several classical models—including long short-term memory (LSTM) networks—showed the RC model consistently achieving lower mean-squared error (MSE) and mean absolute error (MAE), alongside higher $R^{2}$ scores. This work also provided insights into how distinct reservoir initialization strategies influence predictive accuracy, further highlighting the critical role of network dynamics in financial forecasting.
A reservoir is usually represented as a digraph (called the reservoir digraph). In most implementations, reservoir parameters and topological configurations are initialized stochastically. However, such stochastic initialization frequently produces unstable dynamical behavior, necessitating extensive empirical trial-and-error to obtain a well-conditioned network. Therefore, beyond investigating the deployment of reservoirs in prediction tasks, the design of high-quality reservoir architectures has emerged as both a major challenge and a focal point of contemporary research. Gallicchio et al. [12] assessed the impact of constrained reservoir topologies—namely permutation, ring, and chain structures—within Deep Echo State Networks (DeepESNs) on time series forecasting performance. Through a systematic comparison against conventional sparsely random reservoirs and shallow ESNs, demonstrated that embedding structurally sparse topologies in deep reservoir computing constitutes a straightforward yet effective strategy for enhancing temporal modeling accuracy. Cui et al. [19] investigated the interplay between reservoir topology and predictive performance by constructing reservoirs with small-world, scale-free, and hybrid small-world/scale-free structures. Experimental results demonstrate that manually designed reservoir networks deliver superior predictive accuracy and possess a broader spectral radius, while their short-term memory capacity remains comparable to that of traditional models. Moreover, Roeschies et al. [20] proposed an evolutionary algorithm that optimizes reservoir-network architectures for specific problem classes, resulting in substantial gains in predictive performance. Sun et al. [21] introduced a deterministic echo state network architecture featuring a loop reservoir with adjacent feedback connections. This simplified design builds upon the basic loop reservoir structure by incorporating regular adjacent feedback, requiring optimization of only a single parameter. The approach significantly reduces the complexity of ESN implementation while maintaining computational capabilities. These results, together with previous studies, confirm that the structure of a reservoir is pivotal in determining its functionality and overall effectiveness.
2.2 GLMY homology and its applications
GLMY homology represents a significant extension of classical homology theory, specifically developed to analyze directed or asymmetric interactions in digraphs [10] and hypergraphs [22]. This framework has spawned several theoretical developments, such as homotopy theory for digraphs [23], discrete Morse theory on digraphs [24], and path homology theory for multigraphs and quivers [25]. Notable computational advances include the persistent path homology for networks proposed by Chowdhury and Mémoli [11], which offers a novel approach to digraph analysis in topological data analysis, and an efficient algorithm for computing one-dimensional persistent path homology of digraph filtrations by Dey et al. [26]. In contrast to traditional simplicial complexes, GLMY homology offers a unique set of dimensional fingerprints. These fingerprints are instrumental in unearthing the intrinsic structural information hidden within networks, providing a more detailed view of network architecture. This theoretical framework has demonstrated its practical utility across a diverse range of applications, such as modeling intricate biological systems [27], analyzing material properties and structures [28], and understanding complex networks [29, 30]. Persistent GLMY homology represents a further advancement in this area. By tracking how topological features change and evolve across various filtration processes, it allows us to identify and analyze topological patterns in networks [11]. This capability significantly enhances our understanding of network behavior and evolution.
3 Preliminary
In this section, basic concepts of digraphs and persistent homology are introduced.
A digraph is an ordered pair $G=(V,E)$, where $V$ denotes the vertex set and $E\subseteq V\times V$ represents the directed edges. A digraph without loops or multiple edges is called a simple digraph. The GLMY homology (or path homology) provides a topological framework for analyzing simple digraphs [10]. For a finite nonempty vertex set $V$, an elementary $n$-path is defined as a sequence $e_{i_{0}\ldots i_{n}}$ of $n+1$ vertices, where $n\geq 0$. By convention, the set of elementary (1)(-1)-paths is empty. The boundary operator acts on elementary $n$-paths as:
$\partial e_{i_{0}\ldots i_{n}}=\sum_{k=0}^{n}(-1)^{k}e_{i_{0}\ldots\hat{i}_{k}\ldots i_{n}},$
where $\hat{i}_{k}$ indicates the omission of vertex $i_{k}$ from the sequence. An elementary path $e_{i_{0}\ldots i_{n}}$ on a set $V$ is regular if $i_{k-1}\neq i_{k},\forall k=1,\ldots,n$. Otherwise, it is non-regular. Thus, using the vertex set of a simple digraph, we can obtain the set of elementary $n$-paths of $G$, where all elements are regular paths. In the following discussion, unless otherwise noted, simple digraphs are considered.
Let $G=(V,E)$ be a simple digraph. For any integer $n\geq 0$, we define the $\mathbb{K}$-linear space of allowed $n$-paths as:
$\mathcal{A}_{n}(G)=\operatorname{span}\left\{e_{i_{0}\ldots i_{n}}:i_{k}i_{k+1}\in E\text{ for }k=0,\ldots,n-1\right\},$
where $e_{i_{0}\ldots i_{n}}$ denotes an elementary $n$-path. For $n\geq-1$, the subspace of $\partial$-invariant $n$-paths is defined as:
$\Omega_{n}(G)=\left\{v\in\mathcal{A}_{n}:\partial v\in\mathcal{A}_{n-1}\right\},$
with the boundary cases $\Omega_{-1}\cong\mathbb{K}$ and $\Omega_{-2}=\{0\}$. This gives rise to the chain complex:
$\cdots\xrightarrow{\partial}\Omega_{3}\xrightarrow{\partial}\Omega_{2}\xrightarrow{\partial}\Omega_{1}\xrightarrow{\partial}\Omega_{0}\xrightarrow{\partial}\mathbb{K}\xrightarrow{\partial}0.$
The $n$-dimensional GLMY homology group of $G$ is then:
$H_{n}(G)=\operatorname{Ker}(\partial|_{\Omega_{n}})/\operatorname{Im}(\partial|_{\Omega_{n+1}}).$
Elements of $Z_{n}:=\operatorname{Ker}(\partial|_{\Omega_{n}})$ are called $n$-cycles, while elements of $B_{n}:=\operatorname{Im}(\partial|_{\Omega_{n+1}})$ are $n$-boundaries. A representative of a homology generator is any $n$-cycle that is not an $n$-boundary, though such representatives are generally non-unique.
Figure 2: (a) Example of an one-dimensional persistence diagram. (b) An illustration of the confidence set (red area) of (a).
Persistent homology is a tool from computational topology that captures the topological features of space at various scales [31]. It provides a way to quantify the persistence of topological features, capturing the birth and death of topological features in the considered space (such as point clouds, graphs, and digraphs) by constructing a filtration. The persistence diagram (PD) serves as a powerful visualization tool for persistent homology, where each point $(b_{i},d_{i})$ corresponds to a persistence pair representing the birth ($b_{i}$) and death ($d_{i}$) times of a homology generator (Fig. 2). Notice that all points in a PD are located above or on the diagonal. And the value $|d_{i}-b_{i}|$ is called the persistence of the point $(b_{i},d_{i})$, representing the significance of the corresponding topological features. The greater the persistence of a point in the persistence diagram, the more significant the corresponding topological feature is.
A statistical approach for identifying significant points in a persistence diagram is the confidence set [32], which categorizes these points into two classes based on their persistence. As shown in Fig. 2 (b), the confidence set draws two areas on the persistence diagram: the red area and the white area, which contains points with less persistence and greater persistence, respectively. Points in the class that show greater persistence are called significant points, while the other points are called noise points.
The tool of persistence diagram is first employed to illustrate the persistent homology computed from some point clouds, and this framework can also naturally extend to persistent GLMY homology, where we denote the one-dimensional persistence diagram of $\mathbb{H}_{1}$ as $Dgm_{1}(G)$.
4 Methodology
In this section, we first introduce some basic concepts and methods of reservoir computing, as well as some theoretical analysis. In particular, we consider optimization by improving the orthogonality of the reservoir digraph matrix, which enhancing its memory capacity. To achieve this goal, the GLMY theory will be used. Subsequently, we further introduce the persistent GLMY homology of a digraph and the proposed approach for optimizing the reservoirs.
4.1 Reservoir computing and its theoretical analysis
Figure 3: Structure of Reservoir Computing Model
Training the weight matrices in deep learning models is typically computationally expensive. In contrast to conventional neural networks, where the weights of all layers are trained, the reservoir computing (RC) architecture only requires the training of the readout layer’s weights. This characteristic significantly reduces the computational complexity, offering a computational efficiency advantage in certain applications. The general structure of the reservoir computing model is illustrated in Figure 3. It primarily consists of three components: the input layer, the reservoir, and the output layer. The input layer is responsible for encoding the incoming data and feeding it into the reservoir. The reservoir consists of a large, fixed, and recurrent network of neurons that project the input data into a high-dimensional dynamic space. The output layer, which is typically trained, maps the high-dimensional representations from the reservoir to the desired output. Generally, the input and output layers are fully connected layers, while the reservoir is represented by a weight matrix $W_{\text{r}}\in\mathbb{R}^{N\times N}$, where $N$ denotes the size of the reservoir. The weight matrix defines the connections between the neurons within the reservoir, determining the structure and dynamics of the recurrent network. Similarly, the input layer $L_{\text{in}}$ and the output layer $L_{\text{output}}$ are also described by their respective weight matrices. Specifically, the input layer is represented by the weight matrix $W_{\text{in}}\in\mathbb{R}^{N\times L}$, where $L$ is the dimensionality of the input data. The output layer is represented by the weight matrix $W_{\text{out}}\in\mathbb{R}^{M\times N}$, where $M$ is the number of output neurons. These weight matrices govern the flow of information from the input to the reservoir and from the reservoir to the output layer, ensuring the appropriate transformations and mappings of data within the network. Typically, both $W_{\text{in}}$ and $W_{\text{r}}$ are randomly initialized and remain fixed throughout operation, whereas $W_{\text{out}}$ is optimized via gradient descent or ordinary least-squares methods to obtain the optimal parameters. This strategy substantially reduces the training time and computational resources.
At each discrete time step $t+1$, the new reservoir state is jointly determined by the external input arriving at that moment and the internal state carried over from the previous step. Concretely, the update rule is often written as
$\mathbf{x}_{t+1}=(1-\alpha)\mathbf{x}_{t}+\alpha\phi\bigl{(}W_{\text{in}}\mathbf{u}_{t+1}+W_{\text{r}}\mathbf{x}_{t}+\mathbf{b}\bigr{)},$ (1)
where $\mathbf{u}_{t+1}\in\mathbb{R}^{L}$ is the input vector, $\mathbf{x}_{t}\in\mathbb{R}^{N}$ is the previous reservoir state, $W_{\text{in}}\in\mathbb{R}^{N\times L}$ and $W_{\text{r}}\in\mathbb{R}^{N\times N}$ are the fixed input and reservoir weight matrices, $\mathbf{b}$ is an optional bias, and $\phi(\cdot)$ is a point-wise non-linear activation (e.g., $\tanh$). The leak rate $\alpha$ controls the trade-off between retaining historical information and incorporating new information.
More precisely, the activation of each neuron inside the reservoir at time step $t+1$ depends on (i) the subset of input-layer neurons projecting to it and (ii) the neighbouring reservoir neurons to which it is connected. Denoting the $i$-th reservoir neuron by $x_{i,t}$, we have
$x_{i,t+1}=(1-\alpha)x_{i,t}+\alpha\phi\Bigl{(}\sum_{j=1}^{L}W_{\text{in},ij}u_{j,t+1}+\sum_{k=1}^{N}W_{\text{r},ik}x_{k,t}+b_{i}\Bigr{)},$ (2)
where $u_{j,t+1}$ is the $j$-th component of the input vector, $W_{\text{in},\,ij}$ and $W_{\text{r},\,ik}$ are the fixed weights onto neuron $i$, and $b_{i}$ is an optional bias term.
Inspection of the update rules in Eqs. (1)–(2) reveals that, much like a conventional recurrent neural network (RNN)[33], the reservoir maintains a compact representation of both historical and instantaneous information within its state $\mathbf{x}_{t}$. This internal memory allows the model to integrate long-range temporal dependencies when predicting future outcomes. The output at time $t$ is produced by the current reservoir state:
$\mathbf{y}_{t}=W_{\mathrm{out}}\,\mathbf{x}_{t}+\mathbf{c},$
where $\mathbf{y}_{t}\in\mathbb{R}^{M}$ denotes the output vector, $W_{\mathrm{out}}\in\mathbb{R}^{M\times N}$ is the trainable readout weight matrix, and $\mathbf{c}\in\mathbb{R}^{M}$ is an optional bias term. Unlike traditional Recurrent Neural Networks (RNNs) that often struggle with vanishing or exploding gradients when learning long-term dependencies, RC models sidestep this issue by training only a linear readout layer. This simplifies the training process, making RC more stable and efficient in learning the repeating structures characteristic of periodic data.
In our study, we focus on the Echo State Network (ESN), a widely used framework of reservoir computing. Here, an important index that is closely related to the properties of ESN is the (short-term) memory capacity (MC). The memory capacity quantifies the network’s ability to reconstruct past information from its reservoir at the network’s output by computing correlations [34, 35, 36]. It is defined by the following equation:
$\mathrm{MC}=\sum_{k=1}^{k_{\max}}\mathrm{MC}_{k}=\sum_{k=1}^{k_{\max}}\frac{\operatorname{cov}^{2}\left(\mathbf{u}(t-k),\mathbf{y}_{k}(t)\right)}{\operatorname{var}(\mathbf{u}(t))\cdot\operatorname{var}\left(\mathbf{y}_{k}(t)\right)}$ (3)
Here, $\operatorname{cov}$ represents the covariance between the two time series, and $\operatorname{var}$ denotes the variance. $\mathbf{u}(t-k)$ refers to the input presented $k$ steps before the current input, while $\mathbf{y}_{k}(t)=\mathbf{\tilde{u}}(t-k)$ is its reconstructed output from the network using a linear readout. Thus, memory in ESNs is conceptualized as the network’s capacity to retrieve historical information (for various $k$ values) from the reservoir through linear combinations of internal unit activations. The study [36] shows that the memory capacity of ESNs depends on the properties of the structure and parameters of reservoirs, and may also be influenced by the data properties.
A key finding of our study is that introducing additional rings into the reservoir digraph enhances the orthogonality of its adjacency matrix, a property empirically demonstrated to correlate with improved reservoir performance. The orthogonality of reservoir adjacent matrices is closely related to the memory capability of reservoirs, which influences the prediction of reservoirs [37]. Mathematically, orthogonality is a special case of linear independence, which ensures that the matrix columns are not only linearly independent but also mutually orthogonal. Improving the orthogonality of a reservoir matrix reduces overlap and interference between dataset signals stored at different time steps, enabling the network to better distinguish and retrieve distinct inputs. Moreover, orthogonal matrix columns or rows increase the probability of the matrix being full-rank, which is crucial for maximizing memory capacity. Additionally, enhanced orthogonality improves numerical stability, ensuring that computational errors do not accumulate and allowing the network to store and recall information with greater accuracy. In summary, enhancing the orthogonality of reservoir matrices maximizes short-term memory capacity and improves the network’s ability to distinguish and retrieve different input signals [37]. This provides a theoretical base for our optimization method, which focuses on improving the orthogonality of reservoir matrices.
Hence, we aim to increase the orthogonality of the reservoir adjacent matrices to improve their capabilities. Here we use the following measurement to measure the orthogonality of a matrix:
Definition 1.
Suppose $A$ is a matrix. Consider all its column vectors $\{\textbf{a}_{1},\textbf{a}_{2},\cdots,\textbf{a}_{n}\}$, here we define the orthogonality measurement of $A$ is
$OM(A)=\frac{1}{C_{n}^{2}}\sum_{i,j=1;i\neq j}^{n}|m(\textbf{a}_{i},\textbf{a}_{j})|,$
where $C_{n}^{2}=\frac{n(n-1)}{2}$ is the combinatorial number, $m(\textbf{a}_{i},\textbf{a}_{j})=1$ if $\textbf{a}_{i}$ or $\textbf{a}_{j}$ is zero vector, and $m(\textbf{a}_{i},\textbf{a}_{j})=\cos{<\frac{\textbf{a}_{i}}{\|\textbf{a}_{i}\|},\frac{\textbf{a}_{j}}{\|\textbf{a}_{j}\|}>}$ ($\|\cdot\|$ denotes the 2-norm) if both of them are nonzero vectors.
It is easy to see that $OM(A)$ ranges from $[0,1]$, and the closer to zero, the more vectors in $\{\textbf{a}_{1},\textbf{a}_{2},\cdots,\textbf{a}_{n}\}$ are orthogonal to each other, which implies that the orthogonality of this matrix is greater.
Since a reservoir can be represented by a digraph, the problem of improving the orthogonality of its adjacency matrix can also be regarded from the perspective of the reservoir digraph corresponding to the adjacency matrix. To explain this, we first introduce ring, which is a special kind of digraph and is useful for analyzing the properties of reservoirs. A digraph $G=(V,E)$ is a ring if it is connected (as an undigraph), and the in-degree and out-degree of each vertex are 1. We say a digraph has rings if some of its subgraphs are rings. According to the matrix representation of a ring, modifying a reservoir to increase the number of rings in its digraph is equivalent to making its adjacency matrix possess more fundamental cyclic submatrices like:
$\begin{pmatrix}&1&&&\\ &&1&&\\ &&&\ddots&\\ &&&&1\\ 1&&&&\end{pmatrix}$ (4)
or its transpose. For example, in Fig. 4, when changing from (a) to (b), the shape A{A}D{D}B{B}C{C} becomes A{A}D{D}B{B}C{C}, and the corresponding matrix representation changes from $\begin{pmatrix}&1&&1\\ &&1&\\ &&&1\\ &&&\end{pmatrix}$ to $\begin{pmatrix}&1&&\\ &&1&\\ &&&1\\ 1&&&\end{pmatrix}$. Notice that the orthogonality of the adjacency matrix increases from 3 to 4. In fact, for the more general case (Equation 4), since each row and column of this matrix is orthogonal to each other, the more submatrices of this type a matrix has, the greater its orthogonality will be. Combining the previous discussion of the relationship between the orthogonality of the reservoir matrix and the memory capacity of the reservoir, when we make the reservoir matrix (particularly sparse one) have more fundamental cyclic submatrices (which is equivalent to making the digraph of the reservoir have more rings), the greater the memory capacity of the reservoir is, and the more likely it is to exhibit greater predictive ability [37].
Therefore, the core objective of this study is to enhance the orthogonality of the reservoir adjacency matrix. From a graph-theoretic perspective, this goal is equivalent to adding the number of rings in the reservoir digraph—a equivalence that forms the logical starting point for the subsequent methodological design. To achieve this objective, the GLMY homology theory is introduced in this paper as the core analytical tool and operational framework. It is crucial to emphasize that, within the algebraic topological characterization of digraphs, rings are essentially a proper subset of the set of representative cycles contained in the GLMY homology groups. This implies that the representative cycles described by GLMY homology groups include not only rings but also other non-cyclic loop structures.
Based on the above theoretical understanding, the core idea of this study is clarified as follows: to maximize the proportion of rings within the representative cycles of GLMY homology through targeted topological operations. Specifically, it involves implementing structural transformations on those representative cycles that do not currently belong to the ring subset (i.e., non-cyclic loops) to make them become rings, thereby incorporating them into the ring subset. The essence of this process lies in systematically increasing the number of rings in the digraph by optimizing the composition of the representative cycles of GLMY homology, ultimately serving the fundamental goal of improving the orthogonality of the adjacency matrix. The specific implementation method of this transformation process will be introduced in the next subsection.
Figure 4: Illustration of compatibly modifying cycles into rings. (a) A digraph has two rings on the left and right. The middle cycle (red) is not a ring. (b) By changing the direction of $AD$, the middle cycle becomes a ring ABCDABCD compatible with the other rings. (c) The case where the middle cycle cannot be changed into a ring without destroying the other two rings, since either ABAB or DCDC would need to have its direction changed.
4.2 Optimizing reservoir structure by persistent GLMY homology
We first introduce persistent GLMY homology, which extends classical persistent homology theory to the context of weighted digraphs. Let $G=(V,E,w)$ be a weighted digraph with vertex set $V$, edge set $E$, and weight function $w:E\to\mathbb{R}_{+}$. For any threshold $\delta\in\mathbb{R}_{+}$, define the subgraph:
$G^{\delta}=(V,E^{\delta}=\{e\in E:w(e)\leq\delta\}).$
This construction yields a digraph filtration $\{G^{\delta}\hookrightarrow G^{\delta^{\prime}}\}_{\delta\leq\delta^{\prime}}$, where the inclusion maps preserve the vertex set while gradually adding edges in order of increasing weight. The one-dimensional persistent GLMY homology of $G$ is then defined as the persistent vector space:
$\mathbb{H}_{1}=\{H_{1}(G^{\delta})\xrightarrow{i_{\delta,\delta^{\prime}}}H_{1}(G^{\delta^{\prime}})\}_{\delta\leq\delta^{\prime}},$
where $i_{\delta,\delta^{\prime}}$ denotes the induced homomorphism of the inclusion map. Here, each $G^{\delta}$ represents the state of the digraph at filtration parameter $\delta$. This framework provides a systematic way to track the evolution of homology classes as the filtration parameter increases, revealing persistent topological features of the weighted digraph across multiple scales.
In order to find the generators of one-dimensional GLMY homology of a digraph, we first introduce some special structures. A bigon is a sequence of two distinct vertices $a,b\in V$ such that $a\to b,b\to a$. A (boundary) triangle is a sequence of three distinct vertices $a,b,c\in V$ such that $a\to b,b\to c,a\to c$:
b{b}a{a}c{c}
A (boundary) square is a sequence of four distinct vertices $a,b,c,d\in V$ such that $a\to b,b\to c,a\to d,d\to c$:
b{b}c{c}a{a}d{d}
Then, the one-dimensional GLMY homology of simple digraphs is clear because of the following Theorem [26]. This theorem establishes the theoretical foundation for our operation to reservoirs.
Theorem. Let $G=(V,E)$ be a simple digraph. Let $Z_{1}=\operatorname{Ker}\left(\left.\partial\right|_{\Omega_{1}}\right),B_{1}=\operatorname{Im}\left(\left.\partial\right|_{\Omega_{2}}\right)$ , and let $Q$ denote the space generated by all bigons and boundary triangles and boundary squares in $G$. Then, we have $B_{1}=Q$. Hence the one-dimensional GLMY homology group satisfies that $H_{1}=Z_{1}/Q$.
Consequently, any 1-cycle in a digraph that does not form a bigon, boundary triangle, or boundary square serves as a generator of the one-dimensional GLMY homology. In our study, we employ an algorithm that computes the one-dimensional persistent GLMY homology [26] to identify the minimal generators of $H_{1}(G)$ for a given digraph $G$. These minimal generators are called the minimal representative cycles of the $H_{1}(G)$.
Now we introduce our approach for optimizing reservoir structures using GLMY homology to enhance their predictive capabilities. As we have mentioned above, a key distinguishing feature of reservoirs compared to feedforward neural networks is the presence of rings in their architecture, which have been shown to possibly influence their time series prediction performance. Therefore, we aim to increase the number of rings in a given reservoir to optimize its structure. Specifically, when a reservoir is given, we firstly compute the persistent GLMY homology of its digraph and get the minimal representative cycles of one-dimensional GLMY homology groups. Since ring is a special type of minimal representative cycle, we can modified other representative cycles into rings by changing some edges’ directions in a compatible way (this will be introduced later). In other words, these representative cycles are used to modify the structure of the reservoir to increase the number of rings in the reservoir digraph. In our experiments, the reservoir can get better prediction ability on three tested datasets after adding the rings in the reservoir digraph. The flowchart of our method and experimental procedures is shown in Fig. 1.
From a theoretical perspective, the 1-dimensional persistence diagram (1-PD) plays a crucial role in our methodology. Each point in the 1-PD corresponds to a minimal representative cycle in the digraph. Our first step is to check whether each minimal representative cycle forms a closed loop, where edges are connected end-to-end in a sequential and continuous manner, which we define as a "ring". This verification process is essential because only cycles with a proper ring structure can fully contribute to the reservoir’s predictive capabilities. When a minimal representative cycle fails to meet the ring condition, we employ an edge direction modification process. The primary objective of this process is to transform the non-ring representative cycle into a ring while ensuring that all existing rings within the digraph are preserved. For each individual representative cycle, we begin by attempting to modify its edges to achieve the desired ring structure. However, during the edge direction modification, if there are common edges among the current cycle and existing rings, changing the direction of these shared edges may inadvertently destroy these existing rings (see Fig. 4 for an example). In such cases, where altering the current cycle would be incompatible with the preservation of existing rings, we adopt a conservative approach. Instead of forcing the modification and risking the disruption of the existing rings, we abandon the attempt to change the current cycle and proceed to the next representative cycle in the sequence.
To execute this optimization method comprehensively, we apply the above procedure sequentially to all representative cycles corresponding to the points in the 1-PD. Through this iterative process, we gradually and significantly increase the number of rings within the reservoir digraph. As the number of well-formed rings grows, the reservoir’s ability to capture complex patterns improves, leading to enhanced predictive performance. This systematic approach not only leverages the theoretical insights from GLMY homology but also provides a practical and effective solution for reservoir optimization.
The above method can be summarized as the following Algorithm 1.
Algorithm 1 Reservoir Optimization by GLMY Homology
0: A digraph $G=(V,E)$ representing a reservoir.
1: Computing the persistent GLMY homology of $G$ and get the 1-PD.
2: Find the minimal representative cycle for each point in 1-PD. Denote these cycles by $C$.
3:For each cycle in $C$:
4:  If the cycle is already a ring:
5:    Turn to the next cycle.
6:  Else
7:   If the cycle is not a ring but can be modified as a ring without destroying existing rings:
8:     Change some edge directions to make it a ring.
9:   Else
10:    Turn to the next cycle.
11:  EndIf
12:EndFor
12: The optimized digraph $G$ of the reservoir.
Figure 5: Illustration of four datasets, and their Time-delay embedding results in $\mathbb{R}^{3}$ and the one-dimensional persistence diagrams. (a) Result of Mackey-Glass Data. (b) Result of MSO Data. (c) Result of Lorenz System Data. (d) Result of NARMA Data.
5 Experiments and Discussions
In this section, we introduce the datasets used and their topology analysis, and illustrate our experimental results. Our experiments aim to rigorously evaluate prediction performance before and after implementing the optimization process on reservoir digraphs. We will also give some discussions about the proposed method.
In the following subsections, we will use the following designs and parameter settings. In our experimental setup, we assumed that there is at most one edge between every two nodes, and all edges share identical weights for simplicity. The spectral radius  [^2] The spectral radius of a square matrix $A$ is defined as $\rho(A)=\max_{i}|\lambda_{i}|$, where $\{\lambda_{i}\}$ are the eigenvalues of $A$. of the reservoir in our experiments is set as $\lambda_{target}=0.8$ since empirical studies [38] indicate that a spectral radius close to $0.8$ realizes a favourable trade-off between fading memory and nonlinear amplification, thereby yielding best performance for the reservoir. The setting of the spectral radius can be achieved by the following steps: firstly, for a randomly generated reservoir, we assign each edge an identity weight, for example, one, to get the initial reservoir digraph $W_{initial}$. Then we can compute the maximum eigenvalue $\lambda_{max}$ of the reservoir digraph, followed by adjusting the spectral radius of the reservoir to $\lambda_{target}$ using the following formula:
$W_{r}=\lambda_{target}\cdot\frac{W_{initial}}{\lambda_{max}}.$
In the following discussion, we perform 20 tests for each of the initialization method and dataset and take the overall average of each test index, which illustrates the general applicability of our method.
5.1 Datasets and their topology
To validate our approach and compare results from different datasets on the same scale and for the convenience of comparing topological difference, we employ the following datasets [39] with value normalized into range [0.5,0.5][-0.5,0.5]:
  1. 1.
    Mackey-Glass System: This classical chaotic system evaluates time series prediction models through the discrete approximation:
    $y(t+1)=y(t)+\delta\left(a\frac{y(t-\tau/\delta)}{1+y(t-\tau/\delta)^{n}}-by(t)\right),$
    with standard parameters $\delta=0.1$, $a=0.2$, b=0.1b=-0.1, $n=10$, and chaotic regime $\tau=17$. The task involves $k$-step prediction using historical data $u(1,\dots,t)$ to forecast $u(t+k)$. In our experiments, we use the first 5000 steps for training and the following 5000 steps for prediction.
  2. 2.
    Multiple Superimposed Oscillator series: The Multiple Superimposed Oscillator (MSO) series prediction is also an important problem. The core concept behind the MSO is to combine multiple oscillator signals into a single composite indicator by superimposing and normalizing them. It can be described by the following equation:
    $y(t)=\sum_{i=1}^{m}\operatorname{sin}(\alpha_{i}t),$
    where $m$ is the number of summed sine waves. To test ESN, a common way is to set $m=5$ and $\alpha_{1}=0.2,\alpha_{2}=0.311,\alpha_{3}=0.42,\alpha_{4}=0.51,\alpha_{5}=0.63.$
  3. 3.
    Lorenz System: A typical benchmark for chaotic systems:
    $\begin{cases}\dot{x}=-ax(t)+ay(t)\\ \dot{y}=bx(t)-y(t)-x(t)z(t)\\ \dot{z}=x(t)y(t)-cz(t)\end{cases}$
    with $a=10/4$, $b=28$, $c=8/3$. The Runge-Kutta method can generate a sample of length $L$ from the initial conditions $(x_{0},y_{0},t_{0})$ with a step size of 0.01. In our experiments, the $y$-axis is chosen for our prediction task, and we also use the first 5000 steps for training and the following 5000 steps for prediction. The first 1000 input samples were discarded after being fed into the network.
  4. 4.
    NARMA System: This nonlinear auto-regressive moving average system presents a challenging identification problem:
    $y(t+1)=0.3y(t)+$ $0.05y(t)\sum_{i=0}^{9}y(t-i)+$
    $1.5u(t-9)u(t)+0.1,$
    where $u(t)\sim\mathcal{U}[0,0.5]$ (uniform distribution). Initialized with $y(t)=0$ for $t=1,\dots,10$, it tests echo state networks through one-step prediction. In our experiments, we use the first 5000 steps for training and the following 5000 steps for prediction. The first 500 input samples were discarded after being fed into the network.
To analyze the topological differences between datasets, we employ Time-Delay Embedding (TDE) [40] to embed the time series into higher-dimensional space, followed by computing their one-dimensional persistence diagrams (1-PDs) from constructing the alpha filtration [41] of the embedded point clouds. Time-Delay Embedding is a powerful technique used in the analysis of time series data, which transforms a single observed time series into a point cloud in high-dimensional space. This method leverages the concept of phase space reconstruction, allowing for the visualization and analysis of complex time series from limited measurements [40].
Given a scalar time series $\{x_{t}\}_{t=1}^{N}$, where $N$ is the number of observations, the TDE process constructs a set of points in a higher-dimensional space (called the embedding space) using a specified embedding dimension $d$ and a time delay parameter $\tau$. Each point in this new representation is formed as follows:
$\mathbf{X}_{i}=[x_{i},x_{i+\tau},x_{i+2\tau},\ldots,x_{i+(d-1)\tau}]^{T},$
where $i=1,2,\ldots,N-(d-1)\tau$, ensuring that all components of each $\mathbf{X}_{i}$ are within the bounds of the original time series.
The choice of the embedding dimension $d$ and the time delay $\tau$ plays a crucial role in the effectiveness of TDE. The embedding dimension determines the number of past values included in each vector, capturing the dynamics of the underlying system. Meanwhile, the time delay specifies how far apart these values are spaced in time, influencing the independence between the components of the embedded vectors. In our study, by testing a variety of parameter selections, we embed the aforementioned time series into $\mathbb{R}^{3}$ with a delay parameter of $\tau=5$, which gives clear and intuitive illustration of the topological characteristics of the embedded data. This transformation not only enhances the estimation precision of models but also reveals hidden structures and patterns in the data that are not apparent in the raw time series form. The one-dimensional persistence diagram captures the topological features of loops present in the point cloud. When a point in the 1-PD exhibits significantly greater persistence than others, this indicates both the existence of a prominent loop structure in the embedded point cloud and a strong periodic mode in the original time series. This correspondence between persistent homology features and time series periodicity provides valuable insight into the underlying dynamics of the system.
The results are illustrated in Fig. 5, where we illustrate the original time series, the embedded point clouds in $\mathbb{R}^{3}$, and the 1-d persistence diagram with confidence sets. These results reveal distinct topological characteristics: the Mackey-Glass data exhibits a significant 1-dimensional topological feature as its 1-PD has one point locating above the red area, while both the MSO and Lorenz System data display several persistent 1-dimensional features with relatively great persistence, but these points are not regarded as significant. In contrast, the NARMA data is the most obvious one that shows no significant 1-dimensional topological structures. Therefore, these four datasets have different topological features from the perspective of homology.
It is widely observed that Reservoir Computing models often demonstrate superior performance when dealing with datasets exhibiting pronounced periodicity [8, 7]. This enhanced capability stems from RC’s unique architectural features, which are well-suited to capture and exploit the complex temporal dependencies inherent in periodic signals. The fixed, randomly connected “reservoir” acts as a dynamic, high-dimensional feature extractor, mapping input sequences into a rich, non-linear state space. This intrinsic non-linearity and recurrent connectivity enable the reservoir to effectively represent intricate periodic patterns and their underlying dynamics, even when such patterns are embedded within noise or non-stationary components. Furthermore, the inherent memory capacity of the reservoir allows it to retain information about past inputs over extended periods, which is crucial for learning and predicting long-term periodic cycles. According to the operation of TDE, the topology of the embedded data relates to its periodicity, hence influencing the prediction results of the reservoir. This will be shown later in the experimental results.
5.2 Initial reservoir generation and its optimization
In our experiments, we tested three types of initial construction strategies for the reservoir. The first one is constructing the reservoir randomly, which typically results in a sparse reservoir digraph matrix. Additionally, we also test the small-world reservoir constructed with the Watts–Strogatz model [42] and the scale-free reservoir produced by the Barabási–Albert (BA) model [43]. These manually designed network structures have been shown in related studies to provide superior prediction accuracy and a wider spectral radius, which provide a diverse set of benchmarks for our optimization method to validate its effectiveness.
Here, we compute the orthogonality measurement (OM) defined in section.4.1 of each initialization method, comparing their digraph matrices’ OM before and after optimization. Additionally, the number of rings that are added to the digraph is also counted. This evaluation is crucial because research suggests that an increase in the orthogonality of the reservoir matrix can significantly enhance its short-term memory capacity and improve the network’s ability to distinguish and retrieve different input signals. From a topological perspective, increasing the number of rings in the reservoir digraph is equivalent to its adjacency matrix having more fundamental cycle submatrices, which is the key to improving orthogonality.
In this computation, the node number of the reservoir digraph is 500, which means the digraph matrix has the size $500\times 500$. Hence, the orthogonality of each digraph matrix will not be greater than 500. In the Random configuration, we set the sparsity of the reservoir to 0.99, which means that only $1\%$ of the possible edges in the fully connected graph of all nodes, which are randomly selected, actually exist. Due to the presence of a large number of zero elements, the initial orthogonality of the matrix will not be too large. In the small-world configuration, each node was initially connected to its $k=10$ nearest neighbors on a ring lattice, after which every edge was rewired with probability $p=0.1$. In the scale-free configuration, each newly added node established $m=8$ directed links according to preferential attachment.
Table 1 illustrates the comparison of the average orthogonality measurement (OM) and ring addition number from 20 repeated tests of the three different initialization methods, as well as the average optimization running time. For the Random initialization approach, the reservoir’s initial sparsity results in a significant number of zero columns. Our optimization method effectively addresses this by substantially improving the average level of orthogonality, causing the OM to approach zero. Additionally, our approach introduces numerous cycles into the reservoir digraph, strongly demonstrating that the method, based on GLMY homology theory, can indeed optimize the reservoir’s structure. Similar results were observed in the Small-world initialization cases; although the increase in cycles was less than in the Random cases, the average OM was still reduced by approximately half. In the Scale-free approach, the most cycles were added. While its initial OM was already low, our optimization still reduced it by more than half. These results, while showing variations in the original orthogonality and the number of cycles that can be added across different initialization methods, collectively validate our approach’s effectiveness in increasing the orthogonality of the matrices. We can also infer that the orthogonality enhancement will be more significant with larger reservoir matrices.
In the following subsections, we will compute the memory capacity to further validate this observation and provide a more comprehensive performance evaluation. Moreover, we will also provide some related discussions later.
Table 1: Comparison of average OM, ring addition and optimization running time across different initialization methods for $500\times 500$ reservoir digraph matrices.
Random Small-world Scale-free
Original OM 0.3552 0.1866 0.0958
Optimized OM 0.1598 0.0930 0.0438
Added rings 118 76 156
Running time 762.5s 540.5s 347.5s
Figure 6: Average memory capacity (MC) tested on four datasets before and after optimization. Here we compared three initialization approaches.
Figure 7: Illustration of the prediction results on four datasets. (a)–(d) are results of the Mackey-Glass, MSO, Lorenz system, and NARMA dataset, respectively. Figures on the top are prediction results of the initial randomly generated reservoir, with black being the real data and red being the predicted values. It can be seen that after a period of time, the reservoir loses its predictive ability; Figures on the bottom are prediction results of the optimized reservoir, whose predictive ability has been significantly enhanced.
5.3 Memory capacity analysis
In the subsequent experiments, the reservoir digraph’s node number is still set as 500. For the Random initial method, the reservoir sparsity is set to 0.99 (i.e., $1\%$ connectivity), which is a typical setting in practice [16]. This configuration results in a large and sparse reservoir, which effectively leverages the greater predictive power of reservoir computing.
Now we begin introducing the results of the average memory capacity (MC) of reservoirs initialized by the three aforementioned methods, both before and after optimization, which demonstrates the improvement achieved by our approach. Here, the $k_{max}$ in Equation 3 was set to $1.4n$ (where $n$ denotes the reservoir size), following the experimental setup described in [35]. For illustrative purposes, here we test the three initialization methods on the four datasets mentioned before, and compute the average MC of each tested dataset. The results presented in Fig. 6 clearly indicate that our method enhances the MC of the reservoirs. As observed in the figure, although the initial MC values vary across different initialization methods due to their distinct reservoir architectures, our optimization approach exhibits a consistent effect on each method, leading to improvements in MC across the board. Moreover, it can be concluded that the improvement of the MC is related to the type of dataset and the initialization approach. The optimization method consistently and effectively enhances the MC across all tested datasets and initialization schemes. However, the magnitude of this improvement varies significantly. For datasets like Mackey-Glass and MSO, the optimized reservoir shows a dramatic increase in MC, suggesting that the optimization is highly effective for these specific types of time series data. In contrast, for the Lorenz and NARMA datasets, the optimization provides a more modest boost to MC. Furthermore, the choice of initialization method plays a crucial role. For example, the random and small-world initialization consistently yields well improvements of MC values for the Mackey-Glass and MSO tasks, both before and after optimization, highlighting its structural advantage in these scenarios. These results are consistent with our earlier introduction in section 4.1 that MC depends on the properties of the structure and parameters of reservoirs, and may also be influenced by the data properties.
To further visualize these improvements, we showcase some typical experimental results from our study, as depicted in Fig. 7. After optimization, the duration for which the prediction results remain close to the real dataset is significantly extended compared to the pre-optimization state. This extension not only confirms the observable improvement in prediction length but also quantifies its substantial nature, which can be directly attributed to the enhanced memory capacity of the optimized reservoir. Importantly, this phenomenon was consistently observed across lots of the experimental cases, validating its generality.
Therefore, these results demonstrate that after optimization, the reservoirs exhibit enhanced MC, enabling them to better capture the intrinsic features of the datasets. This improvement in feature capture capability directly translates to a significant enhancement in prediction performance. These results illustrate the effectiveness of our optimization method in boosting reservoir computing performance through memory capacity enhancement.
5.4 Prediction performance analysis
Then we compute the average RMSE value tested on four datasets before and after optimization using our approach, illustrating the improvement of performance of reservoirs. The RMSE is computed as
$RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$
where $n$ is the time steps to predict, $y_{i}$ and $\hat{y}_{i}$ represent the predict value and real value of the $i$-step respectively. Table 2 shows the train and test time on the four datasets  [^3] All the experiments were conducted on a server equipped with dual Intel(R) Xeon(R) CPU E5-2678 v3 processors (2.50GHz, 24 cores, 48 threads) and 62 GB of RAM., and Fig. 8 summarizes our experimental results, which shows the tested results on four datasets using three different initial construction. Here, for the Random initialization method, the initial reservoir digraphs are sparse and hence show weak memory capacities and prediction abilities at first, even if the dataset has obvious topological features. However, our experimental results demonstrate that even highly sparse reservoirs can achieve substantially improved prediction performance through adding the number of rings within their digraph architecture. And for the small-world and scale-free methods, since in some studies they are proved to be effective initialization ways of reservoirs [44, 45, 19], we find that they have already achieve relative good performance of prediction on some datasets. But for these cases, we have also observed that after adding the number of rings, if the original reservoir already possessed great prediction capability (such as the small-world and scale-free method on the Lorenz system and NARMA datasets), the modified results can maintain such good capability, i.e., obtaining similar RMSE or lower RMSE. On the contrary, for reservoirs with initially poor prediction capabilities (such as the Mackey-Glass dataset), the optimization process proves to be even more useful, where the modified reservoirs experience a significant enhancement in their prediction power.
Table 2: Comparison of training and testing time (s) across datasets and initialization types.
Dataset Random Small-world Scale-free
Train Test Train Test Train Test
Mackey-Glass 2.1 1.7 3.0 2.7 2.7 2.5
MSO 2.2 1.8 2.7 2.6 2.9 2.6
Lorenz System 2.2 1.8 2.7 2.5 2.6 2.5
NARMA 2.3 1.9 3.0 2.6 2.8 2.8
Moreover, our optimization method yields varying predictive abilities across datasets with distinct topological features. Specifically, the Mackey-Glass dataset demonstrated the most significant improvement in predictive performance. As illustrated by the persistence diagram’s confidence set (Fig. 5), the Mackey-Glass dataset exhibits clear loop features in its embedding space. This aligns perfectly with the conditions under which our method excels, as it leverages these loop structures to enable the optimized reservoir digraph to capture the system’s periodic behaviors effectively. Conversely, the improvement observed with the NARMA dataset was comparatively modest. The NARMA dataset lacks pronounced loop features, which limits our method’s ability to identify strong topological characteristics. Consequently, even after optimization, the performance enhancement remained constrained. The MSO and Lorenz System datasets showed intermediate levels of improvement, falling between the Mackey-Glass and NARMA datasets. While these datasets primarily exhibit noisy topological features, they contain some points with relatively greater persistence compared to the NARMA dataset. Indeed, periodic patterns are discernible in their embedded point clouds. Hence, these two datasets experienced a moderate improvement in the reservoir digraph’s predictive ability after optimization.
Figure 8: Average RMSE tested on four datasets before and after optimizing the reservoir digraph. Here we compared three initialization approaches.
5.5 Discussion
By combining the topological features of the dataset shown in Fig. 5, we found that, under our experimental conditions, a clear pattern emerged regarding the performance of our method: our method demonstrated relatively good performance for datasets with obvious 1-dimensional topological features, namely loops, in the time-delay embedding space. This aligns with the statement in section 4.1, i.e., the reservoir usually performs well on the datasets with pronounced periodicity. In the context of reservoir computing, these 1-d loops represent recurrent structures that can capture temporal dependencies and dynamic behaviors within the data. To predict such kind of dataset (like the Mackey-Glass dataset), successful modeling requires accurate detection and storage of periodic patterns. Reservoir computing architectures are particularly well-suited for this purpose, as their memory capacity enables effective capture of temporal dependencies. This explains why reservoirs with enhanced MC through our optimization method demonstrate good performance on such datasets. When applied to such datasets, our method effectively interacts with these periodic patterns. These features enable the reservoir to better extract and utilize the sequential information, facilitating more accurate representation and forecasting of the underlying system dynamics. Conversely, for datasets with indistinct topological features, the enhancement brought about by our method to the prediction ability of the reservoir digraph was relatively limited. In these cases, the lack of prominent topological structures, such as significant loops, means that the reservoir has fewer obvious patterns to extract when predicting. Without clear topological features, the method’s impact on improving the model’s performance is restricted.
Hence, all these phenomena can be comprehensively understood as the effectiveness of our method being the result of the joint influence of the structural optimization of the reservoir and the characteristics of the dataset itself. On one hand, our method modifies the internal architecture of the reservoir digraph using GLMY homology. On the other hand, the inherent topological features of the dataset act as a foundation, determining the extent to which the method can optimize the model. Only when the method can effectively identify and utilize the topological features of the dataset, as in the case of datasets with distinct one-dimensional loops, can it achieve significant improvements in the prediction performance of the reservoir digraph. Thus, if there is a clear periodic pattern (1-dimensional topological features) in the given time series data, using our optimization method allows the reservoir to generate stronger predictive ability.
In summary, our study not only reveals the performance characteristics of our method under different dataset conditions but also provides valuable insights for future research on integrating topological analysis with model optimization in complex system modeling.
There are some limitations of our method. First, our approach does not capture or modify boundary triangles and boundary squares, as they do not represent generators of GLMY homology despite being cycles. In dense reservoir digraphs, as the connection density is large, boundary triangles can be numerical, making the number of representative cycles be small. But for a sparse reservoir digraph, the boundary cycles are significantly less than the representative cycles, the majority of cycles can still be transformed into rings, thereby enhancing the reservoir’s prediction capability. Hence, our method is more suitable for optimizing sparse reservoirs. And since many rings can share common edges (which means their corresponding fundamental cycle submatrices will share common columns), the increased number of rings is not proportional to improved orthogonality. It is still worthy studying more efficient and reasonable method for improving orthogonality.
Additionally, it is still difficult to give a complete theoretical explanation for why our method can enhance the effect of reservoir, especially when reservoir is a large-scale random sparse matrix, because its properties are difficult to study. Moreover, the computational efficiency also becomes a concern for dense reservoirs, as computing GLMY homology and identifying representative cycles can be time-intensive. Our future research directions will focus on investigating the theoretical relationship between reservoir topology and computational capability, and developing accelerated algorithms for GLMY homology computations.
6 Conclusions
In this study, we proposed a novel approach to optimize the structure of the reservoirs by GLMY homology theory. We begin by computing the one-dimensional GLMY homology of the reservoir digraph to obtain the corresponding minimal representative cycles. These cycles are then utilized to strategically modify the digraph structure in a compatible way, specifically increasing the number of rings within the reservoir. Experimental results demonstrate that this topological optimization leads to improved predictive performance across our tested datasets, and the proposed method performs better on datasets with significant 1-dimensional topological features.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This work was partially supported by National Natural Science Foundation of China under Grant No. 62272406 and Leading Goose R&D Program of Zhejiang under Grant No. 2024C01103.